A chemist wants to mix a 60% saline solution with 20 liters of a 45% saline solution to obtain a 55% saline solution. How much of the 60% solution should she use?
Answer by checkley71(8403)

.6X+.45*20=.55(X+20)
.6X+9=.55X+11
.6X-.55X=11-9
.05X=2
X=2/.05
X=40 LITERS OF 60% SOLUTION IS NEEDED.
PROOF.
.6*40+.45*20=.55(40+20)
24+9=.55*60
33=33